Ok, first of all, I'm not sure how to enter the square root sign, so I will use a v in it's place. Please help me to simplify the following, assuming all variables are non-negative real numbers: 5v1024x^4y^22
Found 2 solutions by stanbon, RAY100:
Answer by stanbon(75887)

5v1024x^4y^22
------------
5th rt(1024x^4y^22)
---
= (1024x^4y^22)^(1/5)
= (2^10*x*4*y^22)^(1/5)
= [(2^10*y^20)^(1/5)]^(1/5)]
= (2^2*y^4)* (x^4y)^(1/5)
= (4y^4)*(x^4*y)^(1/5)


================
Cheers,
Stan H.

Answer by RAY100(1637)

Assuming the problem is:
,
5 * sqrt ( 1024 x^4y^22)
5* sqrt(1024) * sqrt(x^4) * sqrt (y^22)
5 * 32 * x^2 * y^11
160 x^2 y^11

check
let x=3,,,y=4,,,,arbitrarily
original is 5*sqrt ( 1024*3^4*4^22) = 5* sqrt (1.459 x 10^18) =5*1.2 x10^9=6x10^9
answer is 160 *x^2*y^11=160*3^2*4^11=6 x10^9,,,,,,therefore matches